SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Current status of Proto II

[Up] [System Overview] [TNDC Notes] [Operations and IR-2 Data Taking] [Test & Commissioning] [Prototype II] [Mechanical Design] [Safety Systems] [Internal Information] [Electronics] [Software] [Detector Controls]

Last updated: 30 January 2002 by Jignesh Parikh



                                             How Broken Wires were discovered
 

Checking the High Voltage Power Supply

Putting in the new Bertan [17 Jan 02]
Initially, we tried to ramp up the voltages to see if the Bertans were working fine. We found that the Bertan labelled "2020V" was defective. Only one of the two voltages could be turned on. When both "2020V" & "350V" switched on simultaneously, we couldn't ramp up the voltage of anyone. So we changed the Bertan, after which it started working fine.

Testing the NIM power supply [18 Jan 02]
We tried testing the NIM power supply. It was not showing the +12V as it should had showed. Also, while ramping up the voltages, after sometime none of the voltages could be ramped up or down. They would stop working. So we changed the power supply & ensured that it was working fine.

Tripping [18 Jan 02]
We connected the three HV cables to the chamber and turned on the Bertans (no voltage, just the "ON" switches). The "335V" supply started showing current draw up to about 5 uA. On the 1 uA trip range, it would charge up, trip, charge up, trip, etc. On the 10 uA range, it would go up to 0.5 on the front panel meter and would stay there. With the HV cable disconnected from the chamber, the supply would behave perfectly normally. We knew there was something wrong.

Swapping channels of the High Voltage Power  Supply [18 Jan 02]
We then again connected the three HV cables to the chamber and turned on the Bertans. We found that "335V" was still tripping. Also, it was showing some voltage(3V-4V at the output) even when the knob was at zero. So we exchanged the "335V" channel with the 4th unused channel and found that it wasn't tripping.  Also, now it showed zero voltage when the knob was at zero.

Connecting the Power Supply to the chamber [18 Jan 02]
We then  monitered the three voltage test probes from the three bertan channels and the current drawn on "1930V" channel for trips. We increased all the three power supply voltages simultaneously : The power supply "1930V" channel was increased in steps of 100V upto 300V, "870V" in steps of 50V and "335V" in steps of 15V. For "1930V", we tried to put meter range at 1 uA, but it tripped. So we tried putting the meter range at 10 uA, but it was still tripping. So we decided to remove Front & Rear end hardware & trace if there were broken wires.

Hunt  for Broken Wires

Mapping of wires with the design diagram [21 Jan 02]
We got a copy of layout of cables, so that we could mark the cables after checking. We removed the front & rear end hardware. We even removed the Front End Assembly blocks, so that it would be easier for us to trace the cables.

Resistance measurement check for continuity on the same end [22-24 Jan 02]
We started by connecting two field wires on the same end & seeing what resistance it shows. We got some strange results:

  • Field wire 1 in superlayer 1 on south front end showed resistance as 102k  with respect to a reference wire on the same side even though it was not connected and on rear end showed infinite resistance.
  • Field wires 44 & 78 in superlayer 3 on south front end showed resistance as 100k  with respect to a reference wire on the same side even though they were not connected . But when we connected the same wires on rear end they showed infinite resistance.
We concluded that there must be some broken wires.

Continuity Test on front and back end [24-28 Jan 02]
We thought of finding some broken wires externally. So we marked all unconnected wires on the layout. We then started checking each unconnected wire's resistance by connecting the same wire on both ends. We saw some strange observations.

  • We now connected wire 78 on front & rear end and got resistance as 14ohms. When we connected wire 44 on both ends, we got infinite resistance.
  • When we connected sense wires of superlayer 2(south front end) on both ends, they were showing infinite resistance.
We then connected sense wires of superlayer 3(north front end). They looked to be working fine beacuse they were showing some resistance. Field wires of superlayers 2 & 3 were working fine. We didn't know what to do.

Visual Inspection [29 Jan 02]
We then surprisingly came to know that there were small windows on the front & rear end through which you could see inside the proto II. We tried to see through the window if we could see any broken wires. We called Karl Bouldin & Thomas Glanzman to help us. When we were looking through bottom window of the front end, we saw one broken wire but couldn't make out the feedthrough. So Tom, with his digital camera took some pictures of the wire. We saw the pictures on the computer but couldn't locate the feedthrough. So, We went back to take more pictures. Karl finally located the feedthrough with his great eyes, looking through top window of the front end. Tom, tried to take pictures of the feedthrough. We then saw the pictures on the computer & found the exact feedthrough. It was wire 35 of superlayer 4, north front end. We were happy to find the broken wire. But now the bigger problem, what to do next?

List of known Broken Wires [29 Jan 02]

  • Sense Wires
    • [Superlayer 2, South Front End]
      Wire  222          Wire 225         Wire 375
      Wire  319          Wire 372         Wire 278
      Wire  272          Wire 275         Wire 328
      Wire  369          Wire 325         Wire 231
      Wire  322          Wire 228         Wire 378
      Wire 281           Wire 234
     
  • Feild Wires
    • [Superlayer 2, South Front End]    Wire 29
    • [Superlayer 4, North Front End]    Wire 35
  • Zipper Wires
    • [Bottom, South Front End]  Wire 1   Wire 3
Future Plan [31 Jan 02]

Swagato, Mike, Karl & I had a meeting where we discussed the possible options. We agreed upon the fact that replacement of  broken wires would not be advisable till summer. So we stopped monitoring the temperature & water level. We also stopped gas circulation. We are now going to work on setting up the Epics panel for monitoring environment variables in the Proto II.